# Linear analysis of I-C-Love universal relations for neutron stars

### v1

@@ Acknowledgements.
We thank Kent Yagi and Takuya Katagiri for helpful discussions. This work was supported by the National SKA Program of China (2020SKA0120300), the National Natural Science Foundation of China (124B2056), the Beijing Natural Science Foundation (1242018), the Max Planck Partner Group Program funded by the Max Planck Society, and the High-Performance Computing Platform of Peking University. Z.H. is supported by the China Scholarship Council (CSC).
@@ Appendix A Perturbative calculation of $\delta\Lambda/\Lambda$
$+\frac{\partial f}{\partial\rho}\delta\rho+\frac{\partial f}{\partial({\rm d}p/{\rm d}\rho)}\frac{{\rm d}\delta\rho}{{\rm d}p}\,.$
@@
$+\left(\frac{\partial f}{\partial\rho}-\frac{{\rm d}}{{\rm d}p}\frac{\partial f}{\partial({\rm d}p/{\rm d}\rho)}+\frac{\partial f}{\partial y}\frac{\partial f}{\partial({\rm d}p/{\rm d}\rho)}\right)\delta\rho\,,$
@@
$\delta\tilde{y}\equiv\delta y-\frac{\partial f}{\partial({\rm d}p/{\rm d}\rho)}\delta\rho\,,$ (67)
@@
$p=p_{1}\,,\quad\delta\tilde{y}=-\left.\left(\frac{\partial f}{\partial\rho}-\frac{{\rm d}}{{\rm d}p}\frac{\partial f}{\partial({\rm d}p/{\rm d}\rho)}+\frac{\partial f}{\partial y}\frac{\partial f}{\partial({\rm d}p/{\rm d}\rho)}\right)\right|_{p=p_{1}}\,.$ (69)

### v2

@@ Acknowledgements.
We thank Kent Yagi and Takuya Katagiri for helpful discussions, and the anonymous referee for useful comments. This work was supported by the National SKA Program of China (2020SKA0120300), the National Natural Science Foundation of China (124B2056, 12573042), the Beijing Natural Science Foundation (1242018), the Max Planck Partner Group Program funded by the Max Planck Society, and the High-Performance Computing Platform of Peking University. Z.H. is supported by the China Scholarship Council (CSC).
@@ Appendix A Perturbative calculation of $\delta\Lambda/\Lambda$
$+\frac{\partial f}{\partial\rho}\delta\rho+\frac{\partial f}{\partial({\rm d}\rho/{\rm d}p)}\frac{{\rm d}\delta\rho}{{\rm d}p}\,.$
@@
$+\left(\frac{\partial f}{\partial\rho}-\frac{{\rm d}}{{\rm d}p}\frac{\partial f}{\partial({\rm d}\rho/{\rm d}p)}+\frac{\partial f}{\partial y}\frac{\partial f}{\partial({\rm d}\rho/{\rm d}p)}\right)\delta\rho\,$,
@@
$\delta\tilde{y}\equiv\delta y-\frac{\partial f}{\partial({\rm d}\rho/{\rm d}p)}\delta\rho\,$, (75)
@@
$p=p_{1}\,,\quad\delta\tilde{y}=-\left.\left(\frac{\partial f}{\partial\rho}-\frac{{\rm d}}{{\rm d}p}\frac{\partial f}{\partial({\rm d}\rho/{\rm d}p)}+\frac{\partial f}{\partial y}\frac{\partial f}{\partial({\rm d}\rho/{\rm d}p)}\right)\right|_{p=p_{1}}\,.$ (77)
We shall note that, one can directly integral Eq. (73) around $p_{1}$ to obatin the integration initial condition for $\delta y$ in $p=p_{1}^{-}$ that is equivalent to what we have done here. However, one should be careful that $\delta y$ contains a delta function part that also contributes to the jump condition, which is the last term in the above equation.